# Preserve or Modify? Context-Aware Evaluation for Balancing Preservation and Modification in Text-Guided Image Editing

### v2

@@ D.3 Comparison with GPT-4V
Recently, GPT-4V (OpenAI, 2023) has been employed in evaluating various tasks, including text-guided image editing, text-to-image generation, and image quality assessment. Since GPT-4V is one of the best-performing multi-modal large language models, we test the ability of GPT-4V’s effectiveness in evaluating the quality of text-guided edited images. For evaluation, we use the following prompt: “Given a source image and two edited images, you should choose a better edited one based on the source and target text. Source text describes the source image, and target text describes the editing. A well-edited image should preserve the essence of the source image while following the target text.”
As shown in Tab. 12, GPT-4V outperforms AugCLIP in tasks such as EditVal and TEdBench, which involve simple edits like modifying a single object’s attribute. In contrast, our proposed metric, AugCLIP, effectively captures minor differences by augmenting attributes of the source image and target text and shows better performance in other benchmarks with complex scenarios. This finding is consistent with prior research (Zhang et al., 2023), which suggests that GPT-4V struggles to differentiate between images with subtle differences.
@@ D.5 Additional Examples on Problem 2 of Directional CLIP Similarity
To assess directional CLIP similarity’s capability to focus on the image regions modified following the target text rather than peripheral or unchanged regions, we use the relevancy map (Chefer et al., 2021), ${\bm{R}}$. The relevancy map visualizes the transformer’s attention on an image corresponding to a given text depending on their cosine similarity. Specifically, for an image $I\in\mathbb{R}^{h\times w}$ and text $T$, the relevancy map is computed as
@@
Fig. 3(b) and Fig. 14 illustrate the relevancy maps of $\mathrm{CLIP}_{\text{dir}}$ across multiple cases and their improvement achieved by incorporating manually annotated visual descriptions. Unlike $\mathrm{CLIP}_{\text{dir}}$, AugCLIP measures the cosine similarity between the estimated well-edited image and the edited image, rather than between an image and text. As a result, the relevancy map, which requires a direct comparison of the image and text, cannot be applied to AugCLIP.

### v3

@@ D.3 Comparison with GPT-4V
Recently, GPT-4V [22] has been employed in evaluating various tasks, including text-guided image editing, text-to-image generation, and image quality assessment. Since GPT-4V is one of the best-performing multi-modal large language models, we test the ability of GPT-4V’s effectiveness in evaluating the quality of text-guided edited images. For evaluation, we use the following prompt: “Given a source image and two edited images, you should choose a better edited one based on the source and target text. Source text describes the source image, and target text describes the editing. A well-edited image should preserve the essence of the source image while following the target text.”
As shown in Tab. 14, GPT-4V outperforms AugCLIP in tasks such as EditVal and TEdBench, which involve simple edits like modifying a single object’s attribute. In contrast, our proposed metric, AugCLIP, effectively captures minor differences by augmenting attributes of the source image and target text and shows better performance in other benchmarks with complex scenarios. This finding is consistent with prior research [30], which suggests that GPT-4V struggles to differentiate between images with subtle differences.
@@ D.5 Additional Examples on Problem 2 of Directional CLIP Similarity
To assess directional CLIP similarity’s capability to focus on the image regions modified following the target text rather than peripheral or unchanged regions, we use the relevancy map [4], ${\bm{R}}$. The relevancy map visualizes the transformer’s attention on an image corresponding to a given text depending on their cosine similarity. Specifically, for an image $I\in\mathbb{R}^{h\times w}$ and text $T$, the relevancy map is computed as
@@
Fig. 3(b) and Fig. 14 illustrate the relevancy maps of $\mathrm{CLIP}_{\text{dir}}$ across multiple cases and their improvement achieved by incorporating manually annotated visual descriptions. Unlike $\mathrm{CLIP}_{\text{dir}}$, AugCLIP measures the cosine similarity between the estimated well-edited image and the edited image, rather than between an image and text. As a result, the relevancy map, which requires direct comparison of the image and text, cannot be applied to AugCLIP.